OpenStudy (lolox):

Work out the mean of these five expressions : 3n-2,4n,2n+1,n-6,5n-3 Work out the mean of these five expressions : 2n+4,3n-6,4n+1,n-9,7

OpenStudy (tkhunny):

Have you considered adding them and dividing by 5?

OpenStudy (lolox):

But how do i add them when there is n which is unknown @tkhunny

OpenStudy (mathmale):

Example: 3n + 5n = 8n. Add up all of the terms. You don't need a value for n to do that.

OpenStudy (lolox):

Oh okay thank you @mathmale

OpenStudy (tkhunny):

n is just a number. Sometimes it matters what number it is. Oftentimes, it doesn't matter. Just go with it.

OpenStudy (lolox):

I still don't understand what do i do @zenmo

OpenStudy (tkhunny):

Examples: The average of 3 and 5. ==> (3 + 5)/2 = 8/2 = 4 The average of n and 3n. ==> (n + 3n)/2 = 4n/2 = 2n The average of z and 2z and 6z. ==> (z + 2z + 6z)/3 = 9z/3 = 3z The average of 1, 2, 4, 5. ==> (1+2+4+5)/4 = 12/4 = 3 Calculating an average is ALWAYS: 1) Add up your elements 2) Divide that sum by the total number of elements.